OpenStudy (anonymous):

Identify the 19th term of a geometric sequence where a1 = 14 and a9 = 358.80. Round the common ratio and 19th term to the nearest hundredth. a19 ≈ 9,195.53 a19 ≈ 31,035.73 a19 ≈ 13,185.66 a19 ≈ 20,690.49

OpenStudy (callisto):

common ratio \[\sqrt[8]{358.80/14} =?\]

OpenStudy (anonymous):

to be honest idk how to solve that?

OpenStudy (callisto):

First, solve 358.80/14 What do you get?

OpenStudy (anonymous):

25.6285714

OpenStudy (callisto):

Now, use your calculator to solve \[\sqrt[8]{25.62857143}\]

OpenStudy (anonymous):

1.4999975505

OpenStudy (callisto):

Now, r=common ratio a19 = a1 [r^(n-1)] = 14 (1.49999755 ^ (19-1)] = 14 (1.49999755 ^ (18)] = ?

OpenStudy (callisto):

n is the no of term

OpenStudy (anonymous):

20689.878152782416417171686483306764932120668078659492

OpenStudy (anonymous):

would it be a19 ≈ 20,690.49? thats the closest answer

OpenStudy (anonymous):

would it be a19 ≈ 20,690.49? thats the closest answer

OpenStudy (callisto):

Isn't it 20689.88? I don't know :|